What is the force on the charge q at the lower-right-hand corner of the square shown here? (Use the following as necessary for the magnitude: q, a and Eo Assume that the +x-axis is to the right and the ty-axis is up along the page. Enter the direction in degrees counterclockwise from the +x-axis.) magnitude F= direction 45 counterclockwise from the +x-axis
Answer #1

In triangle ABC

using pythagorean theorem

AC = sqrt(AB2 + BC2) = sqrt((a)2 + (a)2) = sqrt(2) a

\theta = tan-1(BC/AB) = tan-1(a/a) = 45

F = force by charge at B and D on charge at C = k q2/a2

F' = force by charge at A on charge at C = k q2/(sqrt(2) a)2 = (0.5)  k q2/a2

net force along the X-direction is given as

Fx = F + F' Cos\theta = (k q2/a2 ) + (0.5) Cos45 k q2/a2 = (1.4) k q2/a2

net force along the Y-direction is given as

Fy = - F - F' Sin\theta = -(k q2/a2 ) - (0.5) Sin45 k q2/a2 = - (1.4) k q2/a2

Net force is given as

F = sqrt(Fx2 + Fy2) = sqrt(2) (1.4) k q2/a2 = (1.98) k q2/a2

direction : \phi = 360 - tan-1(Fy/Fx) = 360 - tan-1(Fy/Fx) = 360 - tan-1(1) = 315